a basketball player shoots the ball with an initial upward velocity of 20ft/s. the ball is 6ft above the floor when it leaves her hands how long will it take for the ball to reach the rim of the basket 10ft above the floor?

Answer by Alan3354(69443)

a basketball player shoots the ball with an initial upward velocity of 20ft/s. the ball is 6ft above the floor when it leaves her hands how long will it take for the ball to reach the rim of the basket 10ft above the floor?
--------------------
The common function used on Earth is
h(t) = -16t^2 + 20t + 6
-16t^2 + 20t + 6 = 10
4t^2 - 5t + 1 = 0
(4t - 1)*(t - 1) = 0
t = 1/4 second going up
t = 1 second falling